2x-4y=7
4x+2y=9
Solve the system by addition. I used the solver but it didn't make sense to me. Can you explain this in more detail please.

Answer by checkley71(8403)

2x-4y=7 & 4x+2y=9 multiply this second equation by 2 we get 8x+4y=18 now add
4x+4y=18
----------
6x=25 or x=25/6 now substituting 25/6 for x in equation we get
2*25/6-4y=7 or 50/6-4y=7 or -4y=7-50/6 or -4y=(42-50)/6 or -4y=-8/6 or
y=(8/6)/4 or y=8/24 or y=1/3
NOTE: The reason for multiplying the equation by 2 is to obtain a +4y term so we can add the two equations & eliminate the y terms (-4y+4y=0). Now we can solve for x.